a - 3 3
------ = -
8 4

Answer by checkley71(8403)

(A-3)/8=3/4 CROSS MULTIPLY
4(A-3)=8*3
4A-12=24
4A=24+12
4A=36
A=36/4
A=9 ANSWER
PROOF
(9-3)/8=3/4
6/8=3/4 CROSS MULTIPLY
24=24